OpenStudy (anonymous):

I've got a little question, don't know if i'm doing it right... 4 = 16 + 9 - 2(12) cosC 0 = 27 - 24cosC 24cosC = 23 23/24 = cosC

OpenStudy (anonymous):

Question :

OpenStudy (anonymous):

the first line is good : 4 = 16 + 9 - 2(12) cosC 0=21-24cosC ..

OpenStudy (anonymous):

oh shoot :p

OpenStudy (anonymous):

just calculation mistake at the second line

OpenStudy (anonymous):

okay nevermind

OpenStudy (anonymous):

thanks :p

OpenStudy (cwrw238):

just a typo